OpenStudy (anonymous):

simplify by reducing the radical . 49√y^21z^14

OpenStudy (amistre64):

turn the exponents into fractions; a sqrt is just a denominator of 2; reduce as needed

OpenStudy (amistre64):

\[\sqrt{y^{21}z^{14}}=y^{21/2}z^{14/2}\]

OpenStudy (anonymous):

the answer is \[49y^{10}z^{7}√y\]

OpenStudy (anonymous):

well the 49 is supposed to look like \[\sqrt[49]{?}\]

OpenStudy (amistre64):

it good to get them into whole numbers as much as possible: 21/2 = 10' 1/2 ; \(y^{10}y^{1/2}\) 14/2 = 7 ; \(x^7\)

OpenStudy (amistre64):

lol ..... then 49 is the denominatir :)

OpenStudy (amistre64):

49rt(.....) is a good way to ascii it

OpenStudy (anonymous):

ooo . so i can have fractions for my exponents?

OpenStudy (amistre64):

\[y^{21/49}\ z^{14/49}\] \[y^{3/7}\ z^{2/7}\] \[\sqrt[7]{y^3\ z^2}\]

OpenStudy (amistre64):

yep; fractional exponents are just a way to write a radical sign

OpenStudy (amistre64):

sqrt = ^1/2 cbrt = ^1/3 4rt = ^1/4 ..... n rt = ^1/n